OpenStudy (anonymous):

A 12 V battery is charged for 20 minutes by connecting it to a source of electromotive force (e.m.f.). The battery is supplied with 7.2 × 104 J of energy in this time. How much charge flows into the battery? A. 5.0 C B 60 C C 100 C D 6000 C answer is D but how?

OpenStudy (anonymous):

E = (1/2) q V, charging of capacitor similar to charging a battery? q = 2 E / V = 2(72000)/12 = 12000 apparently, this analysis is wrong, but I don't see why. E = q V would have given q = E/V = 6000, their answer I guess because the energy storage is due to chemical change rather than capacitance E = q V.